OpenStudy (anonymous):

A conical tent has a radius of 4.1 m and a height of 10.4 m. Doubling which dimension(s) will quadruple the volume of the tent?

OpenStudy (anonymous):

Volume = (1/3) pi r^2 h

OpenStudy (anonymous):

if r is doubled volume quadruples inplace of r,write 2r then \[\left( 2r \right)^2=4r^2\]

OpenStudy (anonymous):

Right.